a Boeing apache longbow military helicopter traveling directly into a strong headwind was able to travel 450 mi in 2.5 h. the return trip, now with a tailwind, took 1 h 40 min. find the speed of the helicopter in calm air and the rate of the wind

Answer by jorel1380(3719)

450/2.5=180 mph
450/1.666=270 mph
180+270/2=225
225-180=45
270-225=45
The speed of the helicopter in calm air is 225 mph. The wind speed is 45 mph..